### v1

# Is My Data in Your Retrieval Database? Membership Inference Attacks Against Retrieval Augmented Generation

Abstract
Retrieval Augmented Generation (RAG) systems have shown great promise in natural language processing. However, their reliance on data stored in a retrieval database, which may contain proprietary or sensitive information, introduces new privacy concerns. Specifically, an attacker may be able to infer whether a certain text passage appears in the retrieval database by observing the outputs of the RAG system, an attack known as a Membership Inference Attack (MIA). Despite the significance of this threat, MIAs against RAG systems have yet remained under-explored.
This study addresses this gap by introducing an efficient and easy-to-use method for conducting MIA against RAG systems. We demonstrate the effectiveness of our attack using two benchmark datasets and multiple generative models, showing that the membership of a document in the retrieval database can be efficiently determined through the creation of an appropriate prompt in both black-box and gray-box settings. Our findings highlight the importance of implementing security countermeasures in deployed RAG systems to protect the privacy and security of retrieval databases.
1 Introduction
Retrieval Augmented Generation (RAG) systems [12, 6] have emerged as a promising approach in natural language processing, gaining significant attention in recent years due to their ability to generate high-quality, up-to-date and contextually relevant responses. These systems combine retrieval and generation components to provide more accurate and informative responses compared to traditional language models. However, like any advanced technology, RAG systems are not immune to vulnerabilities.
While previous research has successfully demonstrated various types of attacks against RAG systems [10, 21, 7, 20], there are still unexplored vulnerabilities in these systems that may pose privacy risks. Specifically, the use of the retrieval database introduces new privacy concerns. Since the retriever component searches a database for relevant passages, an attacker may be able to infer whether a certain text passage appears in the database by observing the outputs of the RAG system. This type of attack is known as a Membership Inference Attack (MIA), and can be used to reveal sensitive information about the contents of the retrieval database.
MIAs were extensively researched in the past in the context of various machine and deep learning systems [15, 3, 2, 18, 9], however to the best of our knowledge the topic of membership inference against RAG systems remains under-explored.
When a MIA is performed against a RAG system it can potentially reveal sensitive or proprietary company information. This may include information about individuals or organizations included in the retrieval database. Furthermore, MIA can be used to prove the unauthorized use of proprietary documents, as part of a legal action [16]. This dual capability makes them a particularly serious form of attack that must be investigated in order to ensure the security and privacy of these systems.
In this study, we introduce a method that is both efficient and easy to use for conducting MIAs against RAG systems, with the objective of determining whether a specific data sample is present in the retrieval database. We assess the effectiveness of our attack using two benchmark datasets that represent scenarios where privacy may be a concern, and employing multiple generative models.
The findings of our study reveal that the membership of a document in the retrieval database can be efficiently determined through the creation of an appropriate prompt, underscoring the importance of implementing security countermeasures in deployed RAG systems.
2 Background
2.1 Retrieval Augmented Generation
Figure 1: Example RAG template for the generation phase of a RAG system. The highlighted placeholders are replaced by the fetched documents from the database and the user prompt, respectively.
Retrieval Augmented Generation is a technique for enriching the knowledge of Large Language Models (LLMs) with external databases without requiring any additional training. This allows easy customization of trained LLMs for specific needs, such as creating AI-based personal assistants or making long textual content (such as a user manual) accessible using simple text queries [4].
Prior to deploying a RAG pipeline, a retrieval database $\mathcal{D}$ must be populated with documents. During this initialization phase, each document is split into chunks, which are mapped into vector representations using a retrieval model $E$, and then stored as an index in the database alongside the original document.
The retrieval model $E$ is specifically designed to learn a mapping from user prompts and documents to a shared vector space, such that prompts are embedded close together with documents containing relevant information to respond to the prompt. This enables efficient searching in the retrieval database, as semantically similar prompts and documents are clustered together in the vector space.
Once deployed, the RAG pipeline typically consists of two main phases: search and generation. In the search phase, $\mathcal{D}$ is queried to find relevant documents that match the user’s query or prompt. When a user prompt is processed by the RAG pipeline, the prompt $p$ is mapped into a vector representation using $E$. Then, $\mathcal{D}$ is searched to find the top-k most similar entries, based on a distance metric calculated using the vector representations (e.g., Euclidean distance). The retrieved entries from $\mathcal{D}$ are organized and provided to the generation phase together with the user prompt.
In the generation phase, a language model $G$ synthesizes the answer based on the retrieved entries from $\mathcal{D}$. The organized information from the search phase is inserted into the RAG template to generate a context $C$. The final system response is obtained by feeding the context $C$, concatenated with the user prompt $p$, into $G$:
$\text{Response}=G(C\parallel p)$ (1)
where $\parallel$ denotes text concatenation.
In practice, it is common to place titles indicating the different areas in the input or additional instructions. For example, placing the sentence ”Answer the question based on the context” after the context and before the user prompt. A full example of a RAG template appears in Figure 1.
2.2 Membership Inference Attacks
Membership inference attacks [15, 9] are a type of privacy threat where an attacker aims to determine whether a specific data record was used in the training set of a machine learning model. This carries significant privacy implications as it can potentially reveal sensitive information about individuals, even if the model does not directly release any personal data.
Formally, an attacker aims to determine the membership of a sample $x$ in the training data of a target model $\mathcal{D}_{m}$, i.e., to check if $x\in\mathcal{D}_{m}$, which is known as sample-level membership inference. Typically, these attacks involve calculating one or more metrics on the target model’s outputs that reflect the probability of the sample being a part of the training set, such as the model outputs’ entropy or log-probabilities [3]. Several metrics may be computed for each sample and then fused together using a machine learning model, known as an attack model, which in turn outputs the probability of a sample being a member of the training set. Additionally, an attacker may also aim to determine the membership of a certain user, i.e., to check if a user’s data is part of the training set, which is known as user-level membership inference [14]. However throughout this paper we will address the membership inference challenge from a sample-level perspective.
In the context of RAG, membership inference can be attributed to either the membership of a sample in the training dataset of the models $E$ or $G$ (described in the previous subsection 2.1), or a document’s membership in the retrieval dataset $\mathcal{D}$. This paper focuses on the latter. Formally, the goal of the attack is to infer the membership of a target document $d$ in the retrieval database $\mathcal{D}$, i.e., to check if $d\in\mathcal{D}$, using only the final output of the RAG system, namely the output of the generative model $G$ conditioned on the fetched context from the retrieval database $\mathcal{D}$.
To the best of our knowledge, this is the first paper to propose such a membership inference attack tailored to RAG systems.
2.3 Threat Model
This paper considers a black-box scenario, in which the attacker has access solely to the user prompt and the resulting generated output from the RAG system. The attacker can modify the user prompt in any manner they deem appropriate; however, they possess no knowledge of the underlying models $E$ or $G$, nor the prompt templates that are being used by these models. Furthermore, the attacker has no information regarding the deployment details, such as the type of retrieval database employed.
In addition to the black-box setting, we also evaluate a supplementary gray-box scenario, in which the attacker has access to the log-probabilities of the generated tokens. Moreover, in this scenario, we assume that the attacker can train the attack model on the target model’s outputs on a subset of the model’s actual training and test datasets [13].
3 Methodology
Figure 2: Attack prompt example for RAG MIA. The highlighted text is the attack-specific part of the prompt, and the rest is taken from the sample for which membership is inferred.
In this section we present our membership inference attack.
The input to the generative model used by the system consists of the RAG tempate into which the retrieved database entries and the user prompt are inserted, as depicted in Figure 1. Since the attacker’s interaction with the system is limited to the user prompt part of the RAG template, the attack must be based on a prompt that achieves both of the following goals:
  • Cause the system to fetch the desired sample from the retrieval database;
  • Cause the system to generate an output that either contains the membership status of the sample, or from which it can be deduced.
The first goal can be easily fulfilled by creating a prompt that has high similarity with the desired sample in the retrieval database. Essentially, any prompt that contains the target sample should be sufficient, as long as there is not too much text added to it. However, the prompt must also cause the generation process to expose the membership status of the sample. Prompting the system with target sample alone will not necessarily achieve this desired result.
To handle this, we designed the following attack prompt format, along with a few additional variations presented in Table 1: ”Does this: ”{Target Sample}” appear in the context? Answer with Yes or No.”. In Figure 2 we present an example of this attack prompt format. Using this prompt we are able to achieve both goals: cause the right sample to be retrieved from the retrieval database, and make the generation phase produce a binary response regarding the sample’s presence in the context. In the black-box setting, we use the model’s answer (Yes/No) alone to deduce the membership status of samples.
As an enhancement to our attack, in cases where the adversary has access to the log-probabilities of the selected tokens, we additionally employ an attack model, similar to the approach described in Section 2.2. In this setup, we apply an ensemble of attack models [13], using as input to the attack both the logits and the class-scaled-logits [3] corresponding to the ”Yes” and ”No” tokens output from the attacked model. The logits are computed by first calculating the exponent of of the log-probability to get a probability estimate $P$ and then applying the logit function.
Since the model only outputs the log-probability of the selected token, for example the ”Yes” token, without the complementary ”No” token, we assign a fixed low probability value of $0.001$ to the complementary token.
Black-Box
Gray-Box
Figure 3: Comparison of different attack prompts. Top: HealthCareMagic, Bottom: Enron
3.1 Experimental Setup
Generative Models
The experiments were conducted on three generative language models:
  • google/flan-ul2 [17], denoted by flan
  • meta-llama/llama-3-8b-instruct [1], denoted by llama
  • mistralai/mistral-7b-instruct-v0-2 [11], denoted by mistral
Datasets
Throughout the evaluation we used two datasets that represent practical scenarios, in which privacy can be critical:
From each dataset, we randomly selected 8,000 samples to be stored in the retrieval database, which we denote as member documents; the remaining 2,000 samples were used as non-member documents in our evaluation.
Embedding Models
The embedding model we used is sentence-transformers/all-minilm-l6-v2, which maps sentences and paragraphs to a 384-dimensional dense vector space.
Retrieval Database
We used a Milvus Lite [19, 8] vector database with $k=4$, Euclidean distance (L2) metric type and HNSW index.
RAG Template
The input to the generative model is built from the user prompt and the context fetched from the retrieval database as a response to the prompt, using the following template:
[…]
In our case, the user prompt was replaced with our special attack prompt.
Attack Prompt
In our evaluation we experimented with 5 different attack prompts, listed in Table 1. Each attack prompt includes a placeholder for a sample, which can be a member or a non-member sample, as shown in Section 3. In the case of the Enron dataset, the sample is the full email body, or its first 1000 characters if it is longer. In HealthCareMagic, the human part of the dialogue is used as the sample.
MIA Attack Details
In both the black-box and the gray-box scenarios we ran our attack using random samples containing 2000 member documents and 2000 non-member documents. In the gray-box case, these documents were split between 40 attack models in the ensemble [13], each model trained 5 times with different train-test splits. In this evaluation, we classified model outputs that failed to include either the ”Yes” or ”No” tokens as non-member documents. This case accounted for approximately 6% of the total outputs. We present the statistics on the amount of such documents in Appendix C.
4 Results
4.1 Impact of Attack Prompt
Table 1: RAG-MIA attack prompts
Attack Prompt
[…]
context?
[…]
the context?
[…]
with Yes or No.
[…]
one word. Answer No if you are not sure.
One important aspect of performing attacks against any generative model is the prompt used to trigger the attack. In the case of RAG, the attack prompt is first used to fetch the relevant documents from the retrieval database, and is then incorporated together with the fetched documents into the RAG template and provided to the generative model. To achieve the best performance, we experimented with five possible attack prompts, described in Table 1. Each one tries to manipulate the RAG system in a slightly different manner to both fetch the relevant documents from the retrieval database, and cause the generative model to output whether the target document was indeed in the retrieval database.
[…]
Unsurprisingly, the results in the black-box setting are inferior to those in the gray-box setting. This is especially the case for the flan model, with an improvement of up to 25% in the gray-box setting. This means that the log-likelihood values for member documents are significantly higher than for non-member documents, indicating a higher confidence of the model in its response.
However, when looking at the llama and mistral models, the average difference is only up to 12%, and in some cases even lower, depending on the prompt.
To further explore this difference, we analyze the percentage of samples that are correctly retrieved from the database for each prompt. We found that over 95% of the member samples are indeed retrieved for both datasets. This is in contrast with the non-member samples, which are retrieved in nearly 0% of the cases. The full results of this analysis can be found in Appendix B.
Thus, we conclude that the flan model is more grounded to the content of its input prompt, and thereby more sure of the presence/absence of a piece of text from it in comparison to the llama and mistral models.
4.2 Results Summary
To quantify the overall MIA risk, we summarize the best results (across prompts) for each combination of model and dataset in Table 2 . We can see that, on average, the attack success rate is highest for flan, in both threat models.
In addition, we observe that the overall risk in the black-box setting is similar between almost all models, with the exception of the HealthCareMagic-llama case. However in the gray-box setting the results vary more.
Our lowest attack AUC-ROC score of 0.74 is quite high compared to previous research on sample-level MIA in language models [5, 14]. Overall, our black-box attack achieves an average AUC-ROC of 0.80 and the gray-box attack achieves an average AUC-ROC of 0.90. The best gray-box attack on flan accomplishes perfect performance. These results highlight the risk associated with deploying RAG based systems without proper defense mechanisms.
Table 2: RAG-MIA results summary
Black-Box
AUC ROC
Gray-Box
AUC ROC
Dataset Model
HealthCareMagic flan 0.80 1.00
llama 0.88 0.96
mistral 0.74 0.83
Enron flan 0.82 0.96
llama 0.79 0.83
mistral 0.78 0.82
5 Conclusion
In this paper we introduced a new membership inference attack against RAG-based systems meant to infer if a specific document was part of the retrieval database or not, a research question previously not explored. Our attack is demonstrated both in black-box and gray-box threat models, and takes advantage of a characteristic of generative models that is usually considered an advantage - context grounding.
Our attack achieves a very high average AUC-ROC of 0.90 and 0.80 in the gray-box and the black-box threat models, respectively, and for some models almost perfect performance. This underscores the need to develop defenses and countermeasures for this attack scenario. We hope that the research community will proceed to explore the risk of membership inference in RAG-based systems and employ the ideas from this paper as a baseline.
6 Acknowledgement
This work was performed as part of the NEMECYS project, which is co-funded by the European Union under grant agreement ID 101094323, by UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee grant numbers 10065802, 10050933 and 10061304, and by the Swiss State Secretariat for Education, Research and Innovation (SERI).
Appendix A Appendices
A Detailed Attack AUC ROC Scores
Table 3 presents the AUC ROC scores for both threat models: black-box and gray-box, and for the different attack prompts.
Table 3: Full RAG-MIA results
Black-Box
AUC ROC
Gray-Box
AUC ROC
Dataset Model
Attack
Prompt
HealthCareMagic flan 0 0.80 0.98
1 0.75 0.97
2 0.78 0.99
3 0.76 0.84
4 0.80 1.00
llama 0 0.82 0.88
1 0.67 0.64
2 0.88 0.96
3 0.78 0.83
4 0.84 0.89
mistral 0 0.71 0.73
1 0.67 0.65
2 0.74 0.83
3 0.71 0.73
4 0.71 0.55
Enron flan 0 0.79 0.93
1 0.68 0.80
2 0.82 0.96
3 0.60 0.62
4 0.80 0.92
llama 0 0.71 0.65
1 0.62 0.54
2 0.79 0.83
3 0.69 0.58
4 0.74 0.74
mistral 0 0.66 0.60
1 0.70 0.68
2 0.78 0.82
3 0.64 0.53
4 0.74 0.69
B Detailed Results of Retrieval Matches
Tables 4 and 5 show the percent of exact matches between the retrieved documents and the member and non-member samples, respectively, for the different attack prompts. We see that the chosen attack prompts do not differ in their influence on the retrieval accuracy.
Table 4: Database retrieval of member documents
Attack
prompt
Equal
count
Total
count
Equal
percent
Dataset
HealthCareMagic 0 1928 2000 96.40
1 1921 2000 96.05
2 1921 2000 96.05
3 1930 2000 96.50
4 1924 2000 96.20
Enron 0 1910 2000 95.50
1 1908 2000 95.40
2 1907 2000 95.35
3 1910 2000 95.50
4 1908 2000 95.40
Table 5: Database retrieval of non-member documents
Attack
prompt
Equal
count
Total
count
Equal
percent
Dataset
HealthCareMagic 0 0 2000 0.00
1 0 2000 0.00
2 0 2000 0.00
3 0 2000 0.00
4 0 2000 0.00
Enron 0 1 2000 0.05
1 1 2000 0.05
2 0 2000 0.00
3 1 2000 0.05
4 0 2000 0.00
C Model outputs without a clear yes/no
answer
The number of model outputs that did not contain either of the ”Yes” or ”No” tokens are shown in Table 6.
Table 6: Model outputs missing yes/no tokens
Missing Total
Percent
missing
Dataset Model
HealthCareMagic flan 923 20000 4.62%
llama 1253 20000 6.27%
mistral 1736 20000 8.68%
Enron flan 1327 20000 6.64%
llama 954 20000 4.77%
mistral 1125 20000 5.63%